OpenStudy (anonymous):

Which of the following equations is the translation 2 units down of the graph of y = |x|? y = |x - 2| y = |x + 2| y = |x| - 2 y = |x| + 2

OpenStudy (anonymous):

down subtract \(2\)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

or just go with C because 92% of the time the answer is C

OpenStudy (igreen):

Lol

OpenStudy (anonymous):

A circle has a diameter with endpoints of (-2, 8) and (6, 4). What is the center of the circle? (8, 12) (4, 12) (2, 6)

OpenStudy (igreen):

C is the correct answer for your first problem.

OpenStudy (anonymous):

thanks both of you

OpenStudy (anonymous):

now the 2nd?

OpenStudy (igreen):

Diameter end points..I assume you must find the point in the middle of them to find the point in the middle of the circle.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

is it 2,6

OpenStudy (igreen):

Yes, that's correct.

OpenStudy (igreen):

As you can see from that graph, the middle of them is (2, 6).

OpenStudy (anonymous):

Which of the following equations is of a parabola with a vertex at (1, -1)? y = (x - 1)2 - 1 y = (x - 1)2 + 1 y = (x + 1)2 - 1 y = (x + 1)2 + 1

OpenStudy (anonymous):

the 2 are squared

OpenStudy (amistre64):

the vertex will zero out the x parts ....

OpenStudy (amistre64):

also, only one of these options works for a point (1,-1)

OpenStudy (anonymous):

A line segment whose endpoints are (4, 2) and (3, 7) is translated to the right 5 units and up one unit. Which of the following matrix sums represents this translation?

OpenStudy (anonymous):

Ask it in a new clean one.